OpenStudy (anonymous):

Two calculations for the area of a circle are shown below. Select the correct calculation explain why you choose this calculation

OpenStudy (anonymous):



OpenStudy (anonymous):

solution 1 A= pie r squared A = 3.14 (30^2) A = 2,826 in ^2

OpenStudy (anonymous):

solution 2 A = pie r squared A = 3.14 (15^2) A = 706.5 in ^2

OpenStudy (anonymous):

@rineheart

OpenStudy (anonymous):

give me a minute to figure it out

OpenStudy (anonymous):

you need to pick the right solution then explain why you choose that one

OpenStudy (anonymous):

its solution two

OpenStudy (anonymous):

explanation

OpenStudy (anonymous):

you multiply the radius sorry lost my train of though

OpenStudy (anonymous):

i need more

OpenStudy (anonymous):

@flcon04 can you help with the explanation

OpenStudy (anonymous):

sorry right now i cant think well enough to help you but i know who can

OpenStudy (anonymous):

@Houdini_Dragon

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@Houdini_Dragon

OpenStudy (anonymous):

@Houdini_Dragon

OpenStudy (anonymous):

@flcon04

OpenStudy (flcon04):

Okay I'm here

OpenStudy (flcon04):

Give me a sec while I figure it out.

OpenStudy (anonymous):

okay

OpenStudy (flcon04):

Okay so, first let me ask what do you think? a or b?

OpenStudy (anonymous):

b

OpenStudy (flcon04):

correct now basically the formula for this is a=pie r^2

OpenStudy (flcon04):

the radius for this circle is 15 because 30 is the diamater so when you half that you get 15

OpenStudy (anonymous):

yes i need to explain why i choose that calculation

OpenStudy (flcon04):

now plug it in and you get A=3.14(15^2)

OpenStudy (flcon04):

step by step 30/2 = 15 (finding the radius) a=3.14(15^2) (plugging in the numbers) now the one which matches with this is B therefore, b is the correct answer!

OpenStudy (anonymous):

thanks for helping her lad